# cryoSPHERE: Single-particle heterogeneous reconstruction from cryo EM

### v1

@@ Abstract
The three-dimensional structure of a protein plays a key role in determining its function. Methods like AlphaFold have revolutionized protein structure prediction based only on the amino-acid sequence. However, proteins often appear in multiple different conformations, and it is highly relevant to resolve the full conformational distribution. Single-particle cryo-electron microscopy (cryo EM) is a powerful tool for capturing a large number of images of a given protein, frequently in different conformations (referred to as particles). The images are, however, very noisy projections of the protein, and traditional methods for cryo EM reconstruction are limited to recovering a single, or a few, conformations. In this paper, we introduce cryoSPHERE, a deep learning method that takes as input a nominal protein structure, e.g. from AlphaFold, learns how to divide it into segments, and how to move these as approximately rigid bodies to fit the different conformations present in the cryo EM dataset. This formulation is shown to provide enough constraints to recover meaningful reconstructions of single protein structures. This is illustrated in three examples where we show consistent improvements over the current state-of-the-art for heterogeneous reconstruction.
Machine Learning, ICML
@@ 1 Introduction
Single-particle cryo-electron microscopy (cryo EM) is a powerful technique for determining the three-dimensional structure of biological macromolecules, including proteins. In a cryo EM experiment, millions of copies of the same protein are first frozen in a thin layer of vitreous ice and then imaged using an electron microscope. This yields a micrograph: a noisy image containing 2D projections of individual proteins. The protein projections are then located on this micrograph and cut out so that an experiment typically yields $10^{4}$ to $10^{7}$ images of size $N_{\textrm{pix}}\times N_{\textrm{pix}}$ of individual proteins, referred to as particles. Our goal is to reconstruct the possible structures of the proteins given these images.
Frequently, proteins are conformationally heterogeneous and each copy represents a different structure. Conventionally, this information has been discarded, and all of the sampled structures were assumed to be in only one or a few conformations (homogeneous reconstruction). Here, we would like to recover all of the structures in a heterogeneous reconstruction.
Structure reconstruction from cryo EM presents a number of challenges. Firstly, each image shows a particle in a different, unknown orientation. Secondly, because of the way the electrons interact with the protein, the spectrum of the images is flipped and reduced. Mathematically, this corresponds to a convolution of each individual image with the Point Spread Function (PSF). Thirdly, the images typically have a very low signal-to-noise ratio (SNR), usually around $0.1$, see e.g. Baxter et al. (2009).
For all these reasons, it is very challenging to perform de novo cryo EM reconstruction. Standard methods, produce electron densities averaged over many, if not all conformations (Scheres, 2012; Punjani et al., 2017), performing discrete heterogeneous reconstruction. More recent methods attempt to extract continuous conformational heterogeneity, e.g. by imposing constraints on the problem through an underlying structure deformed to fit the different conformations present in the dataset, see e.g Rosenbaum et al. (2021); Zhong et al. (2021b); Li et al. (2023). AlphaFold (Jumper et al., 2021) and RosettaFold (Baek et al., 2021) can provide such a structure based on the primary sequence of the protein only. In spite of this strong prior, it is still difficult to recover meaningful conformations. The amount of noise and the fact that we observe only 2D projections creates local minima that are difficult to escape (Zhong et al., 2021b; Rosenbaum et al., 2021), leading to unrealistic conformations.
To remedy this, we root our method in the observation that different conformations can often be explained by large scale movements of domains of the protein (Mardt et al., 2022). Specifically, we develop a variational auto-encoder (VAE) (Kingma & Welling, 2014) that, from a nominal structure and a set of cryo EM images:
@@
These two steps happen concurrently, and the model is end-to-end differentiable. The model is illustrated in Figure 1.
@@
The implementation of the model is available on github  [^1] https://github.com/Gabriel-Ducrocq/VAECryoEM.
@@
2 Background
2.1 Notations and problem formulation
In what follows, $S=\{r_{i}\}_{i=1,\dots,R_{\mathrm{res}}}$ denotes a protein made of a number $R_{\mathrm{res}}\in\mathbb{N}^{\star}$ of residues $r_{i}$. Each residue is itself a set of atoms: $r_{i}=\{a_{k}\}_{k=1,\dots,A_{i}}$ where $A_{i}\in\mathbb{N}^{\star}$ is the number of atoms present in residue $i$. We can equivalently define the structure $S=\{a_{k}\}_{k=1,\dots,A_{\mathrm{total}}}$ in terms of its atoms, where $A_{\mathrm{total}}\in\mathbb{N}^{\star}$ is the total number of atoms in the structure. We denote by $a_{k}=(x_{k},y_{k},z_{k},\eta_{k})^{T}\in\mathbb{R}^{3}\times\{C_{\alpha},C_{\beta},N\}$ the coordinates and atom type of $a_{k}$ for $k\in\{1,\dots,A_{\mathrm{total}}\}$. This means that in this paper, we limit ourselves to the backbone of the protein: we consider only the $C_{\alpha},C_{\beta}$ and $N$ atoms.
@@ 2.1 Notations and problem formulation
The electron density map of a structure $S$, also called a volume, is a function $V_{S}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, where $V_{S}(x)$ represents the probability density function of an electron of $S$ being present, evaluated at $x\in\mathbb{R}^{3}$. That is, if $B\subseteq\mathbb{R}^{3}$, the probability that the structure $S$ has an electron in $B$ is given by $\int_{B}V_{S}(x)\mathrm{d}x$.
@@
Assume we have a set of 2D images $\{I_{i}\}_{i=1}^{N}$ of size $N_{\textrm{pix}}\times N_{\textrm{pix}}$, representing 2D projections of different copies of the same protein in different conformations. Traditionally, the goal of cryo EM heterogeneous reconstruction has been to recover, for each image $i$, the electron density map $V_{i}$ corresponding to the underlying conformation present in image $i$. See Subsection 2.2 for a review these methods. However, following recent works, e.g., Rosenbaum et al. (2021); Zhong et al. (2021b), we aim at recovering, for each image $i$, the underlying structure $S_{i}$ explaining the image. That is, we try to recover the precise position in $\mathbb{R}^{3}$ of each residue.
Since the different conformations of large proteins can often be explained by large scale movements of its domains (Mardt et al., 2022), our method learns to decompose the amino-acid chain of the protein into segments and, for each image $I_{i}$, to rigidly move the learnt segments of a base structure $S_{0}$ to match the conformation present in that image, as explained in Section 1.
AlphaFold (Jumper et al., 2021) and RosettaFold (Baek et al., 2021) can provide such a structure $S_{0}$, based on the amino-acid sequence of the protein at hand. In Section 4, we further fit the AlphaFold predicted structure into a volume recovered by a custom backprojection algorithm provided by Zhong et al. (2020).
2.2 Related work
Two of the most popular methods for cryo EM reconstruction, which are not based on deep learning, are RELION (Scheres, 2012) and cryoSPARC (Punjani et al., 2017). Both methods perform volume reconstruction, hypothesize that $k$ conformations are present in the dataset and perform maximum a posteriori estimation over the $k$ density maps, thus performing discrete heterogeneous reconstruction. Both of these algorithms operate in Fourier space using an expectation-maximization algorithm (Dempster et al., 1977) and are non-amortized: the poses are learnt for each image, and adding new images to the dataset would require a new run of the algorithm. Other approaches perform continuous heterogeneous reconstruction. For example, 3DVA (Punjani & Fleet, 2021b) uses a probabilistic principal component analysis model to learn a latent space while E2GMM (Chen & Ludtke, 2021) parameterizes the volume as a mixture of $N$ Gaussian p.d.f and learns their mean conditionally on an input image.
Another class of methods involve deep learning and typically performs continuous heterogeneous reconstruction using a VAE architecture. Some of them attempt to reconstruct a density map, while others target the atomic structures – or a backbone – directly. For example, cryoDRGN (Zhong et al., 2020, 2021a) and CryoAI (Levy et al., 2022) use a VAE acting on Fourier space to learn a latent space and a mapping that associates a 3D density map with each latent variable. They perform non-amortized and amortized inference over the poses, respectively. Other methods are defined in the image space, e.g. 3DFlex (Punjani & Fleet, 2021a) and cryoPoseNet (Nashed et al., 2021). They both perform non-amortized inference over the poses. These methods either learn, for a given image $I_{i}$, $\{V_{i}(x_{k})\}$ the values at a set of $N_{\textrm{pix}}^{3}$ fixed 3D coordinates $\{x_{k}\}$, representing the volume on a grid (explicit parameterization), or an actual function $\hat{V}_{i}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ in the form of a neural network that can be queried at chosen coordinates (implicit parameterization). The implicit parameterization requires smaller networks but needs an expensive $N_{\textrm{pix}}^{2}$ and $N_{\textrm{pix}}^{3}$ evaluation for image and volume generation, respectively.
Other deep learning methods attempt to directly reconstruct structures instead of volumes and share a common process: starting from a plausible base structure, obtained with e.g. AlphaFold (Jumper et al., 2021), for each image, they move each residue of the base structure to fit the conformation present in that specific image. These methods differ on how they parameterize the structure and in the prior they impose on the deformed structure or the motion of the residues. For example AtomVAE (Rosenbaum et al., 2021) considers only residues and penalizes the distances between two subsequent residues that deviate too much from an expected value. CryoFold (Zhong et al., 2021b) considers the residue centers and their side-chain and also imposes a loss on the distances between subsequent residues and the distances between the residue centers and their side-chain. Unfortunately, due to the high level of noise and the fact that we observe only projections of the structures, these ”per-residue transformation” methods tend to be stuck in local minima, yielding unrealistic conformations unless the base structure is taken from the distribution of conformations present in the images (Zhong et al., 2021b), limiting their applicability to real datasets. Even though AtomVAE (Rosenbaum et al., 2021) could roughly approximate the distribution of states of the protein, it was not able to recover the conformation given a specific image. In addition, to our knowledge, the code for these two methods is not available and non-trivial to reimplement. This prevents a comparison of our results with AtomVAE and CryoFold. To reduce the bias that the base structure brings, (Schwab et al., 2023) recently proposed to fit pseudo-atoms in a consensus map with a neural network directly. Similar to our work, several other methods constrain the atomic model to rigid body motions. For example (Chen et al., 2023), starting from a base structure $S_{\mathrm{big}}$ fitted in the consensus map (called the big GMM), learn a representation structure $S_{\mathrm{small}}$ of this consensus map using a smaller number using the method described in (Schwab et al., 2023). They subsequently learn a translation on each residue $S_{\mathrm{small}}$, that contribute to translate the residues of $S_{\mathrm{big}}$ weighted by how close they are from the residues of $S_{\mathrm{small}}$. This can be thought of as our segmentation GMM, except that it takes places in the 3D space and only the centers of the Gausian modes can be learnt, while ours is fitted on the residue indexes and all of the GMM parameters are learnable. In addition, they do not use this segmentation to perform rigid body motions. Instead, in a subsequent step, the user is in charge of identifying the rigidly moving segments of the protein by applying one or several masks to the protein. This is in contrast to our method where the segmentation used to perform rigid body motion is learnt by the network trying to move the domains to fit the images. Finally, this method involves a sequence of optimization problem, with potentially different architecture, while ours is end-to-end differentiable. Using the method in (Schwab et al., 2023), Chen et al. (2024) develop a focused refinement on patches of the GMM representation of the protein. These patches are learnt using k-means on the location of residues and do not depend on the different conformations of the data set. This in contrast to cryoSPHERE which learns the segments of the protein are tightly linked to the change of conformation. Concurrently to our work, Li et al. (2023) developed CryoSTAR which acts on each residue and enforces local rigidity of the motion of the protein by imposing a similarity loss between the base structure and the deformed structure. The interested reader can see Donnat et al. (2022) for an in-depth review of deep learning methods for cryo EM reconstruction.
The reconstruction methods relying on an atomic model, such as e.g cryoSTAR, DynaMight or cryoSPHERE offer the possibility to the user to provide prior information via this atomic model. They also offer the possibility of deforming the protein according to chemical force fields. This is not the case of the methods performing volume reconstruction without such an atomic model.
@@ 3 Method – cryoSPHERE
In this section, we present our method for single-particle heterogeneous reconstruction, denoted cryoSPHERE. The method focuses on structure instead of volume reconstruction. It differs from the previous works along this line in the way the movements of the residues are constrained: instead of deforming the base structure on a residue level and then imposing a loss on the reconstructed structure, it moves parts of the protein approximately rigidly by construction.
We use a typical VAE architecture, see Figure 1 for a depiction of how the information flows in our network. We map each image to a latent variable by a stochastic encoder. The latent variable is then passed to a decoder which outputs a transformation (rotation and translation) per segment. Based on these transformations and segment decomposition, the underlying structure $S_{0}$ is deformed and then turned into a volume. Finally, the volume is projected into a 2D image, according to its pose that can be estimated using existing software, e.g. (Scheres, 2012; Punjani et al., 2017). Finally, this image is convolved with the PSF to give the reconstructed image that is compared to the input image. This completes the forward pass. In the backward pass, the parameters of the encoder, decoder and Gaussian mixture are updated with a gradient descent based on the evidence lower bound (see Section 3.2). In the following, we describe details of our network: the encoder and decoder, the segment decomposition, how to turn a structure into a 2D projection of its electron density, and an efficient way to compare the true and predicted images.
@@ 3.1 Image formation model
$V_{S}(r):=\sum_{a\in S}\dfrac{1}{\sqrt{2\pi}\sigma}\exp\left({-\dfrac{||r-a||^{2}}{2\sigma^{2}}}\right)$ (1)
where $r\in\mathbb{R}^{3}$ and $\sigma$ is a positive real number. In this definition the protein’s electron density is approximated as the sum of a Gaussian p.d.f centered on each atom. We use a fixed value of $\sigma=1$, following Rosenbaum et al. (2021).
From these density maps, we then compute an image projection $I\in\mathbb{R}^{N_{\textrm{pix}}\times N_{\textrm{pix}}}$ as:
@@
where $(r_{x},r_{y})\in\mathbb{R}^{2}$ are the coordinates of a pixel, $r_{z}\in\mathbb{R}$ is the coordinate along the $z$ axis, $R\in SO(3)$ is a rotation matrix and $t\in\mathbb{R}^{3}$ is a translation vector. The abuse of notation $RS+t$ means that every atom of $S$ is rotated according to $R$ and then translated according to $t$. The image is finally convolved with the point spread function (PSF) $g$, which in Fourier space is the contrast transfer function (CTF), see Vulović et al. (2013).
In practice, we avoid approximating the integral in Equation (2). The density map, being a sum of Gaussian p.d.f, allows for an exact calculation by integrating over the last marginal. This significantly reduces the computing time.
Figure 4: Spike dataset. Left: Predicted angles. Middle left: FSC curves for the open and close conformations. Middle: Close conformation. Middle right: Open conformation. Right: Definition of the angles.
@@ 3.3 Segment decomposition
We fix a maximum number of segments $N_{\textrm{segm}}\in\{1,\dots,R_{\mathrm{res}}\}$ and we represent the decomposition of the protein by a stochastic matrix $G\in\mathbb{R}^{R_{\mathrm{res}}\times N_{\textrm{segm}}}$. The rows of $G$ represent ”how much of each residue belongs to each segment”, and our objective is to ensure that each residue mostly belongs to one segment, that is:
@@
We also aim for the segments to respect the sequential structure of the amino acid chain, and the model to be end-to-end differentiable. Without end-to-end differentiability, we could not apply the reparameterization trick and we would have to resort on Monte-Carlo estimation of the gradient of the segments, which has a higher variance, see e.g (Mohamed et al., 2019).
To meet these criteria, we fit a Gaussian mixture model (GMM) with $N_{\textrm{segm}}$ components on the real line supporting the residue indices, that is, on $\mathbb{R}$. Each component $m$ has a mean $\mu_{m}$, standard deviation $\sigma_{m}$ and a softmax weight $\alpha_{m}$. The $\{\alpha_{m}\}$ are passed into a softmax to obtain the weights $\{\pi_{m}\}$ of the GMM, ensuring they are positive and summing to one. We define the probability that a residue $i$ belongs to segment $m$ as the probability that the index $i$ belongs to mode $m$, annealed by a temperature $\tau\in\mathbb{R}^{+\star}$:
@@
In this ”soft” decomposition of the protein, each residue can belong to more than one segment, allowing for smooth deformations. In addition, the differentiable architecture is amenable to gradient descent methods, and a well chosen $\tau$ can approximate a ”hard” decomposition of the protein. We set $\tau=20$ in the experiment section. In our experience, this segmentation procedure is very robust to different initialization and converge in only a few epochs.

### v2

@@ Abstract
The three-dimensional structure of proteins plays a crucial role in determining their function. Protein structure prediction methods, like AlphaFold, offer rapid access to a protein’s structure. However, large protein complexes cannot be reliably predicted, and proteins are dynamic, making it important to resolve their full conformational distribution. Single-particle cryo-electron microscopy (cryo-EM) is a powerful tool for determining the structures of large protein complexes. Importantly, the numerous images of a given protein contain underutilized information about conformational heterogeneity. These images are very noisy projections of the protein, and traditional methods for cryo-EM reconstruction are limited to recovering only one or a few consensus conformations. In this paper, we introduce cryoSPHERE, which is a deep learning method that uses a nominal protein structure (e.g., from AlphaFold) as input, learns how to divide it into segments, and moves these segments as approximately rigid bodies to fit the different conformations present in the cryo-EM dataset. This approach provides enough constraints to enable meaningful reconstructions of single protein structural ensembles. We demonstrate this with two synthetic datasets featuring varying levels of noise, as well as two real dataset. We show that cryoSPHERE is very resilient to the high levels of noise typically encountered in experiments, where we see consistent improvements over the current state-of-the-art for heterogeneous reconstruction.
@@ 1 Introduction
Single-particle cryo-electron microscopy (cryo-EM) is a powerful technique for determining the three-dimensional structure of biological macromolecules, including proteins. In a cryo-EM experiment, millions of copies of the same protein are first frozen in a thin layer of vitreous ice and then imaged using an electron microscope. This yields a micrograph: a noisy image containing 2D projections of individual proteins. The protein projections are then located on this micrograph and cut out so that an experiment typically yields $10^{4}$ to $10^{7}$ images of size $N_{\textrm{pix}}\times N_{\textrm{pix}}$ of individual proteins, referred to as particles. Our goal is to reconstruct the possible structures of the proteins given these images. Frequently, proteins are conformationally heterogeneous and each copy represents a different structure. Conventionally, this information has been discarded, and all of the sampled structures were assumed to be in only one or a few conformations (homogeneous reconstruction). Here, we would like to recover all of the structures in a heterogeneous reconstruction.
Structure reconstruction from cryo-EM presents a number of challenges. First, each image shows a particle in a different, unknown orientation. Second, because of the way the electrons interact with the protein, the spectrum of the images is flipped and reduced. Mathematically, this corresponds to a convolution of each individual image with the Point Spread Function (PSF). Third, the images typically have a very low signal-to-noise ratio (SNR). For these reasons, it is very challenging to perform de novo cryo-EM reconstruction. Standard methods, produce electron densities averaged over many, if not all conformations (Scheres, 2012; Punjani et al., 2017), performing discrete heterogeneous reconstruction. More recent methods attempt to extract continuous conformational heterogeneity, e.g., by imposing constraints on the problem through an underlying structure deformed to fit the different conformations present in the dataset, see e.g. Rosenbaum et al. (2021); Zhong et al. (2021b); Li et al. (2023). AlphaFold (Jumper et al., 2021) and RosettaFold (Baek et al., 2021) can provide such a structure based on the primary sequence of the protein only. In spite of this strong prior, it is still difficult to recover meaningful conformations. The amount of noise and the fact that we observe only 2D projections creates local minima that are difficult to escape (Zhong et al., 2021b; Rosenbaum et al., 2021), leading to unrealistic conformations.
To remedy this, we root our method in the observation that different conformations can often be explained by large scale movements of domains of the protein (Mardt et al., 2022). Specifically, we develop a variational auto-encoder (VAE) (Kingma & Welling, 2014) that, from a nominal structure and a set of cryo-EM images:
@@
These two steps happen concurrently, and the model is end-to-end differentiable. The model is illustrated in Figure 1. The implementation of the model is available on github  [^1] https://github.com/Gabriel-Ducrocq/cryoSPHERE.
@@
2 Notations and problem formulation
@@
In what follows, we consider only the $C_{\alpha}$ atoms of the protein. A protein made of a number $R_{\mathrm{res}}\in\mathbb{N}^{\star}$ of residues $r_{i}$ is denoted $S=\{r_{i}\}_{i=1}^{R_{\mathrm{res}}}$, where the coordinates of residue $i$ are the coordinates of its $C_{\alpha}$ atom. The electron density map of a structure $S$, also called a volume, is a function $V_{S}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, where $V_{S}(x)$ is proportional to the probability density function of an electron of $S$ being present in an infinitesimal region around $x\in\mathbb{R}^{3}$. That is, the expected number of electrons in $B\subseteq\mathbb{R}^{3}$ is proportional to $\int_{B}V_{S}(x)\mathrm{d}x$.
@@
Assume we have a set of 2D images $\{I_{i}\}_{i=1}^{N}$ of size $N_{\textrm{pix}}\times N_{\textrm{pix}}$, representing 2D projections of different copies of the same protein in different conformations. Traditionally, the goal of cryo-EM heterogeneous reconstruction has been to recover, for each image $i$, the electron density map $V_{i}$ corresponding to the underlying conformation present in image $i$; see Section 4 for a review of these methods. However, following recent works, e.g., Rosenbaum et al. (2021); Zhong et al. (2021b), we aim at recovering, for each image $i$, the underlying structure $S_{i}$ explaining the image. That is, we try to recover the precise position in $\mathbb{R}^{3}$ of each residue.
@@ 3 Method – cryoSPHERE
In this section, we present our method for single-particle heterogeneous reconstruction, denoted cryoSPHERE. The method focuses on structure instead of volume reconstruction. It differs from the previous (Rosenbaum et al., 2021) and concurrent (Li et al., 2023) works along this line in the way the movements of the residues are constrained: instead of deforming the base structure on a residue level and then imposing a loss on the reconstructed structure, our method learns to decompose the amino-acid chain of the protein into segments and, for each image $I_{i}$, to rigidly move the learnt segments of a base structure $S_{0}$ to match the conformation present in that image. This is motivated by the fact that different conformations of large proteins often can be explained by large scale movements of its domains (Mardt et al., 2022).
The base structure $S_{0}$ can be obtained using methods like AlphaFold (Jumper et al., 2021) and RosettaFold (Baek et al., 2021), based on the amino-acid sequence of the protein. In Section 5, we further fit the AlphaFold predicted structure into a volume recovered by a custom backprojection algorithm provided by Zhong et al. (2020).
We use a type of VAE architecture, see Figure 1. We map each image to a latent variable by a stochastic encoder, which is then decoded to a rigid body transformation per segment. Based on these transformations and the segment decomposition, the underlying structure $S_{0}$ is deformed, posed and turned into a volume that is used to create a projected image. This image is then compared to the input image. After that, the backward pass updates the parameters of the encoder, decoder and Gaussian mixture. We now describe the details of our model.
@@ 3.1 Image formation model
$V_{S}(r):=\sum_{a\in S}A_{a}\exp\left({-\dfrac{||r-a||^{2}}{2\sigma^{2}}}\right)$ (1)
where $A_{a}$ is the average number of electrons per atom in residue $a$, $r\in\mathbb{R}^{3}$ and $\sigma=2$ by default. Hence, the protein’s electron density is approximated as the sum of Gaussian kernels centered on its $C_{\alpha}$ atoms. From these density maps, we then compute an image projection $I\in\mathbb{R}^{N_{\textrm{pix}}\times N_{\textrm{pix}}}$ as:
@@
where $(r_{x},r_{y})\in\mathbb{R}^{2}$ are the coordinates of a pixel, $r_{z}\in\mathbb{R}$ is the coordinate along the $z$ axis, $R\in SO(3)$ is a rotation matrix and $t\in\mathbb{R}^{3}$ is a translation vector. The abuse of notation $RS+t$ means that every atom of $S$ is rotated according to $R$ and then translated according to $t$. The image is finally convolved with the point spread function (PSF) $g$, which in Fourier space is the contrast transfer function (CTF), see Vulović et al. (2013). Note that the integral can be computed exactly for our choice of approximating the density map as a sum of Gaussian kernels, which significantly reduces the computing time.
@@ 3.3 Segment decomposition
To handle the often very low SNR encountered in cryo-EM data, we regularize the transformation of the structure produced by the decoder by restricting it to transforming whole segments of the protein. We fix a maximum number of segments $N_{\textrm{segm}}\in\{1,\dots,R_{\mathrm{res}}\}$ and we represent the decomposition of the protein by a stochastic matrix $G\in\mathbb{R}^{R_{\mathrm{res}}\times N_{\textrm{segm}}}$. The rows of $G$ represent ”how much of each residue belongs to each segment”, and our objective is to ensure that each residue primarily belongs to one segment, that is:
@@
We also aim for the segments to respect the sequential structure of the amino acid chain, and the model to be end-to-end differentiable. Without end-to-end differentiability, we could not apply the reparameterization trick and we would have to resort to Monte Carlo estimation of the gradient of the segments, which has a higher variance, see e.g. Mohamed et al. (2019).
To meet these criteria, we fit a Gaussian mixture model (GMM) with $N_{\textrm{segm}}$ components on the real line supporting the residue indices. Each component $m$ has a mean $\mu_{m}$, standard deviation $\sigma_{m}$ and a logit weight $\alpha_{m}$. The $\{\alpha_{m}\}$ are passed into a softmax to obtain the weights $\{\pi_{m}\}$ of the GMM, ensuring they are positive and summing to one. We further anneal the Gaussian components by a temperature $\tau>0$, and define the probability that a residue $i$ belongs to segment $m$ as:
@@
In this ”soft” decomposition of the protein, each residue can belong to more than one segment, allowing for smooth deformations. In addition, the differentiable architecture is amenable to gradient descent methods, and a well chosen $\tau$ can approximate a ”hard” decomposition of the protein. We set $\tau=20$ in the experiment section. In our experience, this segmentation procedure is very robust to different initialization and converges in only a few epochs.